A ball weighing 1kg, suspended on a spring, oscillates harmoniously with an amplitude of 10 cm and a period T = 3s. Determine the kinetic energy when the ball is in equilibrium
1
Expert's answer
2021-03-18T14:04:01-0400

"\text{according to the law of conservation of energy:}"

"E_k+E_p= const"

"E_p=\frac{kx^2}{2}"

"x\text{ is the distance from the equilibrium point}"

"T=2\pi\sqrt{\frac{m}{k}}"

"k=(\frac{2\pi}{T})^2m=(\frac{2\pi}{3})^2*1\approx2.09"

"E{p}\text{ maximum if }x=0.1"

"E_{pmax}=\frac{kx^2}{2}=\frac{2.09*0.1^2}{2}=0.01"

"\text{if }x=0\ E{kmax}=E_{pmax}"

"E_k(0) = 0.01J"

Answer: 0.01 J kinetic energy when the ball is in equilibrium